OpenStudy (anonymous):

The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2).

OpenStudy (texaschic101):

y = 1/2x + 6 can you tell me the slope of this line ?

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position. So in y = 1/2x + 6, the number in the m position is 1/2, and that is your slope. Parallel lines will have the same slope. y = mx + b slope(m) = 1/2 (0,-2) x = 0 and y = -2 now sub and solve for b -2 = 1/2(0) + b -2 = b Therefore, your parallel equation is : y = 1/2x - 2